A Bibliography of Papers in Lecture Notes in Computer Science (1990)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org, beebe@ieee.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

29 July 2017 Version 1.10

Title word cross-reference

2n [13]. Step_{k,m} [67]. N [99, 13]. P [71].

-Bit [13]. -Cascades [67]. -Dimensional [99].

Abstract [62, 21, 98, 107, 104, 54, 31, 100, 17, 10, 75, 85, 120, 86, 128, 111, 68, 127, 69, 36, 115, 41]. Abuse [75]. Abuse-Free [75]. Access [102, 64]. Active [32]. Addition [110]. Adolescence [1]. Age [61]. Agent [125]. Algebraic [63]. Algebraic-Code [63]. Algorithm [109, 88]. Algorithms [13, 28]. Alive [6]. Almost [60]. Alternative [19]. and/or [44]. Antique [94]. Any [10, 115].

Application [43]. Applications [121, 123, 17, 116, 111, 89]. Architecture [106]. Argued [21]. Atkin [59]. Attack [56, 132, 88]. Attacks [32]. Authentication [92, 46, 65, 2, 48, 78, 12]. Authentications [116]. Authenticity [6]. Author [73, 134]. Avalanche [9].

Based [92, 40, 104, 10, 5, 99, 109, 127]. Batch [90]. Be [51, 21]. Binary [70]. Bit [76, 86, 13]. Bits [123, 24]. Block [13, 115]. Bounded} [21]. Bounds [12]. Boxes [129]. Break [37]. Bronze [61]. Brujin [52].

Call [132, 28]. Can [51, 21, 83]. Can’t [51]. Card [23]. Cards [68, 95]. Cartesian [46]. Cascades [67]. Case [61]. Cash [30, 116]. Certain [70]. Certified [94]. Chain [110].

Checks [31, 30]. Checksum [88]. Children [133]. Chip [24]. Chosen [88]. Cipher [58, 13]. Ciphers [115]. Classification [98]. Code [63]. Codes [12]. Cohen [88]. Collision [43, 111]. Commitment [76, 86]. Communication [103]. Commutative [16]. Complexity [51, 118, 57, 70, 54, 50, 55]. Computation [124]. Computational [22]. Computationally [69]. Computations [125]. Computer [6, 130]. Concrete [118]. Conditionally [75]. Conditions [70]. Confidential [64]. Congruential [87]. Consistency [89]. Constructing [41]. Construction [115]. Context [75]. Continued [55]. Control [102, 44]. Controlled [123]. Convert [130]. Correlation [56]. Counting [9]. Crete [61]. Criteria [53]. Criterion [9]. Cross [99]. Cross-Product [99]. Cryptanalysis [40, 67, 39, 36, 89]. Cryptel [25]. Crypto [81]. Crypto-Engine [81]. Cryptographers [69]. Cryptographic [53, 88, 80, 28]. Cryptographically [125]. Cryptography [77, 1]. Cryptoperiods [14]. Cryptosystem [72, 4, 114, 132]. Cryptosystems [91, 71, 75, 101, 63, 7]. Curve [72, 91]. Curves [40].

Data [131]. DC [33]. Deciphering [61]. Deficiencies [18]. Defined [52]. Definition [99]. Describe [51]. Design [71, 112]. Detection [33]. Diffusing [14]. Digital [92, 97, 94, 103]. Dimensional [99]. Dining [69]. Direct [37]. Disclose [17]. Disclosure [123]. Disco [69]. Diskette [81]. Disposable [116]. Disrupters [33]. Distributed [27, 130]. Distribution [10, 103, 105]. Distributions [85]. Divergence [12]. Diversity [60]. Divertible [16]. Domains [100]. Double [125]. Double-Agent [125]. Dynamic [99].

Easy [129, 43, 111]. Efficiency [57]. Efficient [117, 31, 128, 68, 95, 127, 23].

Electronic [31, 25, 35, 116]. Elliptic [72, 91]. Encryption [40, 78, 106]. End [106]. End-to-End [106]. Engine [81]. Enigma [74]. European [28]. Evaluation [28]. Every [50]. Everything [21]. Exchange [126, 57, 104, 65, 5]. Existence [76]. Existing [25]. Explain [133]. Exponent [11]. Exponentiation [108]. Exponents [36]. Extended [62, 21, 98, 104, 54, 31, 100, 10, 75, 85, 120, 86, 43, 64, 127, 22, 115, 41].

Facing [29]. Factoring [35]. Fast [72, 56, 60, 109]. Faster [62]. Faulty [124, 122]. FEAL [132]. FEAL- [132]. Feed [56]. Feed-Forward [56]. Feedback [54, 83, 66]. Feedforward [52]. Fiat [19]. Fields [104]. Filling [40]. Filtered [56]. Find [129]. Five [23]. Flexible [102]. Forward [56]. Fractions [7, 55]. Free [75]. Front [59]. Full [65]. Function [10]. Functions [52, 112, 9, 53, 113, 13].

Gamal [4]. General [15, 75].

Generalization [4]. Generate [83]. Generation [49, 60]. Generator [42]. Generators [87]. German [74]. Given [83]. Giving [18]. Good [129, 50]. Gradual [123]. Group [8].

Half [122]. Hardware [107]. Hash [112, 113, 13]. Hash-Functions [13]. Heuristics [110]. High [63]. Higher [9, 109]. Hints [18]. Homophonic [38]. Huang [88]. Hypotheses [115].

Ideal [98, 45]. Identification [117, 68, 95, 127]. Identity [5]. Identity-Based [5]. If [51]. Implementation [91, 37]. Implementations [107]. Impossibility [41]. Improve [3]. Improved [58]. Index [73, 134]. Infinite [100]. Information [63, 38]. Information-Theoretic [38].

Informational [12]. Integrity [28]. Interactive [117, 16, 121, 92]. Invited [29, 81, 77, 107, 1, 61, 79, 34, 74, 78, 44, 80].

Keep [6]. Kerberos [78]. Kernels [127]. Key [126, 58, 57, 104, 10, 1, 65, 2, 48, 5, 63, 7, 4, 103, 105]. Key-Exchange [5]. Keying [74]. Keys [102]. Keystream [50]. Knowledge [92, 118, 18, 21, 20, 15, 76, 119, 120, 16, 116, 133, 64, 22].

Large [8]. Later [79]. LCT [89]. Legal [29]. Line [97]. Line/Off [97]. Linear [70, 54, 99, 50, 55, 89, 61]. linearity [11]. Local [84].

Machine [39]. Mail [25, 35]. Majorities [125]. Majority [124]. Making [75, 23]. Mapping [34]. Master [102]. Match [23]. Match-Making [23]. Maximal [60]. Medical [64]. Message [92]. Minimal [48]. Minimum [120]. Minorities [125]. MIXes [37]. Mobile [103]. Modified [88, 39]. Modular [108, 109]. Modular-Multiplication [109]. Moduli [60]. Multiparty [124, 122, 125]. Multiple [58]. Multiplication [109].

Navy [74]. Network [6, 78]. Networks [77, 8]. News [59]. No [47, 65]. Non [121, 92, 11]. Non-interactive [121, 92]. Non-linearity [11]. Nonlinearity [53]. Nonlinearly [56]. Normal [77]. NP [21]. Number [21].

Oblivious [121]. Offline [31]. On-Line [97]. On-Line/Off-Line [97]. One [10, 113]. One-Way [10]. Online [30]. Open [27]. Optimally [41]. Order [9]. Oriented [8].

Paper [94]. Paradigms [92]. Parallel [49]. Password [79]. Off-Line [97]. or [44]. Perfect [84, 21]. Periodic [82]. Permutation [71]. Permutations [11, 41].

Permuted [127]. Personal [81]. PGM [114]. Point [50, 26]. Power [22]. Practical [18, 25, 8, 80]. Predict [87]. Prepositioned [44]. Previously [65]. Primality [62]. Primitives [28]. Principle [112]. Private [63]. Private-Key [63]. Problem [125]. Problems [77, 80]. Procedures [64]. Processing [27, 130]. Processor [24]. Processors [122]. Product [99]. Profile [50]. Profiles [70, 55]. Progress [131]. Proofs [92, 117, 118, 18, 76, 119, 120, 16, 22]. Properties [114]. Protection [25, 128, 80]. Protocol [33, 8, 5, 19, 103]. Protocols [122, 126, 133]. Provably [115]. Prover [117]. Providing [77]. Pseudo [51, 84, 86, 42]. Pseudo-random [51, 84]. Pseudo-randomness [86]. Pseudorandom [85, 41]. Public [1, 2, 7, 4]. Public-Key [1, 2, 7].

Quadratic [104, 82].

Race [28]. Radix [109]. Random [123, 34, 16, 42, 51, 84]. Randomness [84, 86]. Rates [63]. Real [104]. Recipient [69, 32]. Records [64]. Recurring [49]. Recursive [108]. Reducibility [16]. Register [56, 83]. Registers [54, 66]. Relying [115]. Requirements [29]. Residues [108]. Resource [120]. Results [111, 41]. RIPE [28]. Rotor [39]. Rounds [21, 119]. RSA [107, 90, 60, 37, 24, 36]. RSA-Implementation [37]. RSA-Moduli [60].

S [129]. S-Boxes [129]. Satisfiability [23]. Satisfying [9]. Say [47]. Scheme [58, 15, 2, 99, 128, 127, 80, 96]. Schemes [117, 3, 98, 45, 123, 76, 46, 44]. Schnorr [42]. Scripts [61]. SDNS [106]. Search [43, 111]. Secrecy [48]. Secret [126, 98, 45, 100, 44, 36]. Secrets [17, 65]. Secure [125, 75, 65, 2, 60, 69, 115]. Security [77, 79, 131, 42, 27, 26]. Self [16].

Self-Reducibility [16]. Sender [69, 32]. Sequence [51, 83]. Sequences [51, 70, 82, 52, 56, 49, 84, 50]. Serviceability [69]. Shamir [19]. Shared [65, 44, 96]. Sharing [98, 45, 100]. Shift [56, 83, 66]. Shift-Register [56]. Short [36]. Shortest [83]. Signature [29, 3, 94, 96]. Signatures [92, 93, 97, 68, 95]. Simple [14]. Single [24]. Smart [81, 68, 95]. Software [128]. Some [45, 70]. Sorting [20]. Space [40, 99]. Space-Filling [40]. Spans [82]. Sparse [85]. spite [32]. Spymasters [125]. Standardisation [131]. Starting [50, 26]. Statistics [34]. Strict [9]. Structure [126]. Substitution [38]. Subtitle [94]. Summary [43, 64, 22]. Sums [108]. Survey [107]. Symmetric [13]. System [57, 104, 25, 10]. Systems [48, 103, 105].

Technical [26]. Technique [14]. Technology [29]. Ten [79]. Test [59, 89]. Testing [62]. Text [88]. Their [123, 116]. Theoretic [38]. Threshold [101, 99]. Token [81]. Tolerating [122]. Transfer [121]. Treatment [38]. Trick} [23]. Two [117, 119]. Type [71].

Unclassified [77]. Unconditional [48, 69, 32]. Unconditionally [125, 75]. Undeniable [93]. Universal [81]. UNIX [79]. Unproved [115]. Untraceability [69, 32]. Untraceable [116]. Use [77, 7, 78]. User [81]. Using [117, 18, 108, 86, 13].

Varying [66]. Verifiable [17]. Verification [96]. Video [40]. Viruses [130]. Voting [58].

Way [10, 113].

Years [79].

Zero [92, 118, 18, 21, 20, 15, 76, 119, 120, 16, 116, 133, 64, 22]. Zero-Knowledge [118, 18, 21, 20, 15, 76, 120, 116, 133, 64, 22].

References

Diffie:1990:APK

[1] Whitfield Diffie. The adolescence of public-key cryptography (invited). Lecture Notes in Computer Science, 434: 2–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                             tronic). URL http://link.springer-
                                             ny.com/link/service/series/0558/
                                             bibs/0434/04340002.htm;   http://
                                             link.springer-ny.com/link/service/
                                             series/0558/papers/0434/04340002.
                                             pdf.

Galil:1990:SPK

[2] Zvi Galil, Stuart Haber, and Moti Yung. A secure public-key authentication scheme. Lecture Notes in Computer Science, 434:3–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                             tronic). URL http://link.springer-
                                             ny.com/link/service/series/0558/
                                             bibs/0434/04340003.htm;   http://
                                             link.springer-ny.com/link/service/
                                             series/0558/papers/0434/04340003.
                                             pdf.

Brassard:1990:HIS

[3] Gilles Brassard. How to improve signature schemes. Lecture Notes in Computer Science, 434:16–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                             http://link.springer-ny.com/link/
                                             service/series/0558/bibs/0434/04340016.
                                             htm;   http://link.springer-ny.
                                             com/link/service/series/0558/papers/
                                             0434/04340016.pdf.

Jaburek:1990:GGP

[4] W. J. Jaburek. A generalization of El Gamal’s public key cryptosystem. Lecture Notes in Computer Science, 434:23–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340023.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340023.
pdf.

Gunther:1990:IBK

[5] Christoph G. Günther. An identitybased key-exchange protocol. Lecture Notes in Computer Science, 434:29– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340029.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340029.
pdf.

Bauspiess:1990:HKA

[6] Fritz Bauspieß and Hans-Joachim Knobloch. How to keep authenticity alive in a computer network. Lecture Notes in Computer Science, 434:38– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340038.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340038.
pdf.

Isselhorst:1990:UFP

[7] Hartmut Isselhorst. The use of fractions in public-key cryptosystems. Lecture Notes in Computer Science, 434: 47–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340047.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340047.
pdf.

Frankel:1990:PPL

[8] Yair Frankel. A practical protocol for large group oriented networks. Lecture Notes in Computer Science, 434: 56–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340056.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340056.
pdf.

Lloyd:1990:CFS

[9] Sheelagh Lloyd. Counting functions satisfying a higher order strict avalanche criterion. Lecture Notes in Computer Science, 434:63–??, 1990. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/0434/04340063.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0434/04340063.pdf.

Davida:1990:KDS

[10] George I. Davida, Yvo Desmedt, and René Peralta. A key distribution system based on any one-way function (extended abstract). Lecture Notes in Computer Science, 434:75– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340075.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340075.
pdf.

Pieprzyk:1990:NLE

[11] Józef P. Pieprzyk. Non-linearity of exponent permutations. Lecture Notes in Computer Science, 434:80– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340080.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340080.
pdf.

Sgarro:1990:IDB

[12] Andrea Sgarro. Informational divergence bounds for authentication codes. Lecture Notes in Computer Science, 434:93–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340093.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340093.
pdf.

Quisquater:1990:BHF

[13] Jean-Jacques Quisquater and Marc Girault. 2n-bit hash-functions using nbit symmetric block cipher algorithms. Lecture Notes in Computer Science, 434:102–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/0434/04340102.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340102.
                                        pdf.

Mjolsnes:1990:STD

[14] Stig F. Mjølsnes. A simple technique for diffusing cryptoperiods. Lecture Notes in Computer Science, 434: 110–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340110.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340110.
                                        pdf.

Burmester:1990:GZK

[15] Mike V. D. Burmester, Yvo Desmedt, Fred Piper, and Michael Walker. A general zero-knowledge scheme. Lecture Notes in Computer Science, 434: 122–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340122.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340122.
                                        pdf.

Okamoto:1990:DZKa

[16] Tatsuaki Okamoto and Kazuo Ohta. Divertible zero knowledge interactive proofs and commutative random self-reducibility. Lecture Notes in Computer Science, 434:134–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340134.htm;   http://

link.springer-ny.com/link/service/
series/0558/papers/0434/04340134.
pdf.

Crepeau:1990:VDS

[17] Claude Crépeau. Verifiable disclose for secrets and applications (abstract). Lecture Notes in Computer Science, 434:150–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340150.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340150.
pdf.

Boyar:1990:PZK

[18] Joan Boyar, Katalin Friedl, and Carsten Lund. Practical zero-knowledge proofs: Giving hints and using deficiencies. Lecture Notes in Computer Science, 434:155–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340155.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340155.
pdf.

Stern:1990:AFS

[19] Jacques Stern. An alternative to the Fiat–Shamir protocol. Lecture Notes in Computer Science, 434:173– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340173.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340173.
pdf.

Brassard:1990:SZK

[20] Gilles Brassard and Claude Crépeau. Sorting out zero-knowledge. Lecture Notes in Computer Science, 434:181– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340181.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340181.
                                        pdf.

Brassard:1990:ENC

[21] Gilles Brassard, Claude Crépeau, and Moti Yung. Everything in NP can be argued in Perfect zero-knowledge in a Bounded number of rounds (extended abstract). Lecture Notes in Computer Science, 434:192–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340192.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340192.
                                        pdf.

Yung:1990:ZKP

[22] Moti Yung. Zero-knowledge proofs of computational power (extended summary). Lecture Notes in Computer Science, 434:196–??, 1990. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/series/0558/bibs/0434/04340196.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0434/04340196.pdf.

denBoer:1990:MEM

[23] Bert den Boer. More efficient match-making and satisfiability: The Five Card Trick. Lecture Notes in Computer Science, 434:208–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340208.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340208.
pdf.

Vandemeulebroecke:1990:SCB

[24] André Vandemeulebroecke, Etienne Vanzieleghem, Tony Denayer, and Paul G. A. Jespers. A single chip 1024 bits RSA processor. Lecture Notes in Computer Science, 434:219– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340219.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340219.
pdf.

Cnudde:1990:CPP

[25] Hedwig Cnudde. Cryptel — the practical protection of an existing electronic mail system. Lecture Notes in Computer Science, 434:237–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340237.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340237.
pdf.

VanAuseloos:1990:TSS

[26] Jan Van Auseloos. Technical security: The starting point. Lecture Notes in Computer Science, 434:243– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340243.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340243.
                                        pdf.

Siuda:1990:SOD

[27] Charles Siuda. Security in open distributed processing. Lecture Notes in Computer Science, 434:249–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340249.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340249.
                                        pdf.

Vandewalle:1990:ECC

[28] Joos Vandewalle, David Chaum, Walter Fumy, Cees J. A. Jansen, Peter Landrock, and G. Roelofsen. A European call for cryptographic algorithms: RIPE: Race Integrity Primitives Evaluation. Lecture Notes in Computer Science, 434:267–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340267.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340267.
                                        pdf.

Antoine:1990:LRF

[29] Mireille Antoine, Jean-François Brakeland, Marc Eloy, and Yves Poullet. Legal requirements facing new signature technology (invited). Lecture Notes in Computer Science, 434:273– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340273.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340273.
pdf.

Chaum:1990:OCC

[30] David Chaum. Online cash checks. Lecture Notes in Computer Science, 434:288–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340288.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340288.
pdf.

Chaum:1990:EOE

[31] David Chaum, Bert den Boer, Eugène van Heyst, Stig F. Mjølsnes, and Adri Steenbeek. Efficient offline electronic checks (extended abstract). Lecture Notes in Computer Science, 434:294– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340294.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340294.
pdf.

Waidner:1990:USR

[32] Michael Waidner. Unconditional sender and recipient untraceability in spite of active attacks. Lecture Notes in Computer Science, 434:302– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340302.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340302.
pdf.

Bos:1990:DDD

[33] Jurjen N. E. Bos and Bert den Boer. Detection of disrupters in the DC protocol. Lecture Notes in Computer Science, 434:320–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340320.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340320.
pdf.

Flajolet:1990:RMS

[34] Philippe Flajolet and Andrew M. Odlyzko. Random mapping statistics (invited). Lecture Notes in Computer Science, 434:329–354, 1990. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/0434/04340329.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0434/04340329.pdf;   http://www.
research.att.com/~amo/doc/arch/
random.mappings.pdf;   http://www.
research.att.com/~amo/doc/arch/
random.mappings.ps;   http://www.

research.att.com/~amo/doc/arch/
random.mappings.tex.

Lenstra:1990:FEM

[35] Arjen K. Lenstra and Mark S. Manasse. Factoring by electronic mail. Lecture Notes in Computer Science, 434:355–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340355.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340355.
pdf.

Wiener:1990:CSR

[36] Michael J. Wiener. Cryptanalysis of short RSA secret exponents (abstract). Lecture Notes in Computer Science, 434:372–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340372.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340372.
pdf.

Pfitzmann:1990:HBD

[37] Birgit Pfitzmann and Andreas Pfitzmann. How to break the direct RSAimplementation of MIXes. Lecture Notes in Computer Science, 434:373– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340373.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340373.
pdf.

Jendal:1990:ITT

[38] Hakon N. Jendal, Yves J. B. Kuhn, and James L. Massey. An informationtheoretic treatment of homophonic substitution. Lecture Notes in Computer Science, 434:382–??, 1990. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/series/0558/bibs/0434/04340382.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0434/04340382.pdf.

Wichmann:1990:CMR

[39] Peer Wichmann. Cryptanalysis of a modified rotor machine. Lecture Notes in Computer Science, 434:395– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340395.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340395.
                                        pdf.

Bertilsson:1990:CVE

[40] Michael Bertilsson, Ernest F. Brickell, and Ingemar Ingemarsson. Cryptanalysis of video encryption based on space-filling curves. Lecture Notes in Computer Science, 434:403–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340403.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340403.
                                        pdf.

Zheng:1990:IOR

[41] Yuliang Zheng, Tsutomu Matsumoto, and Hideki Imai. Impossibility and optimally results on constructing pseudorandom permutations (extended abstract). Lecture Notes in Computer Science, 434:412–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340412.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340412.
pdf.

Rueppel:1990:SSP

[42] Rainer A. Rueppel. On the security of Schnorr’s pseudo random generator. Lecture Notes in Computer Science, 434:423–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340423.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340423.
pdf.

Quisquater:1990:HECa

[43] Jean-Jacques Quisquater and Jean-Paul Delescaille. How easy is collision search? application to DES (extended summary). Lecture Notes in Computer Science, 434:429–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340429.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340429.
pdf.

Simmons:1990:PSS

[44] Gustavus J. Simmons. Prepositioned shared secret and/or shared control schemes (invited). Lecture Notes in Computer Science, 434:436– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340436.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340436.
                                        pdf.

Brickell:1990:SIS

[45] Ernest F. Brickell. Some ideal secret sharing schemes. Lecture Notes in Computer Science, 434:468–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340468.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340468.
                                        pdf.

DeSoete:1990:CAS

[46] Marijke De Soete, Klaus Vedder, and Michael Walker. Cartesian authentication schemes. Lecture Notes in Computer Science, 434:476–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340476.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340476.
                                        pdf.

Beutelspacher:1990:HSN

[47] Albrecht Beutelspacher. How to say “no”. Lecture Notes in Computer Science, 434:491–??, 1990. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/0434/04340491.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0434/04340491.pdf.

Godlewski:1990:KMA

[48] Philippe Godlewski and Chris Mitchell. Key minimal authentication systems for unconditional secrecy. Lecture Notes in Computer Science, 434:497– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340497.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340497.
pdf.

Gunther:1990:PGR

[49] Christoph G. Günther. Parallel generation of recurring sequences. Lecture Notes in Computer Science, 434: 503–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340503.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340503.
pdf.

Niederreiter:1990:KSG

[50] Harald Niederreiter. Keystream sequences with a good linear complexity profile for every starting point.

Lecture Notes in Computer Science, 434:523–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340523.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340523.
pdf.

Beth:1990:CPR

[51] Thomas Beth and Zong-Duo Dai. On the complexity of pseudo-random sequences — or: If you can describe a sequence it can’t be random. Lecture Notes in Computer Science, 434: 533–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340533.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340533.
pdf.

Dai:1990:FFD

[52] Zong-Duo Dai and Kencheng Zeng. Feedforward functions defined by de brujin sequences. Lecture Notes in Computer Science, 434:544–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340544.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340544.
pdf.

Meier:1990:NCC

[53] Willi Meier and Othmar Staffelbach. Nonlinearity criteria for cryptographic functions. Lecture Notes in Computer Science, 434:549–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340549.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340549.
pdf.

Chan:1990:LCF

[54] Agnes Hui Chan, Mark Goresky, and Andrew Klapper. On the linear complexity of feedback registers (extended abstract). Lecture Notes in Computer Science, 434:563– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340563.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340563.
pdf.

Wang:1990:LCP

[55] Muzhong Wang. Linear complexity profiles and continued fractions. Lecture Notes in Computer Science, 434: 571–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340571.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340571.
pdf.

Forre:1990:FCA

[56] Réjane Forré. A fast correlation attack on nonlinearly feed-forward filtered shift-register sequences. Lecture Notes in Computer Science, 434:

586–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340586.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340586.
                                        pdf.

Buchmann:1990:CEN

[57] Johannes A. Buchmann, Stephen Düllmann, and Hugh C. Williams. On the complexity and efficiency of a new key exchange system. Lecture Notes in Computer Science, 434: 597–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340597.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340597.
                                        pdf.

Boyd:1990:NMK

[58] Colin Boyd. A new multiple key cipher and an improved voting scheme. Lecture Notes in Computer Science, 434:617–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340617.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340617.
                                        pdf.

Morain:1990:ATN

[59] François Morain. Atkin’s test: News from the front. Lecture Notes in Computer Science, 434:626–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340626.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340626.
pdf.

Maurer:1990:FGS

[60] Ueli M. Maurer. Fast generation of secure RSA-moduli with almost maximal diversity. Lecture Notes in Computer Science, 434:636–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340636.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340636.
pdf.

Duhoux:1990:DBA

[61] Yves Duhoux. Deciphering Bronze Age scripts of Crete — the case of Linear A (invited). Lecture Notes in Computer Science, 434:649–650, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340649.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340649.
pdf.

Bosma:1990:FPT

[62] Wieb Bosma and Marc Paul van der Hulst. Faster primality testing (extended abstract). Lecture Notes in Computer Science, 434:652–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340652.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340652.
                                        pdf.

Hwang:1990:PKA

[63] Tzonelih Hwang and T. R. N. Rao. Private-key algebraic-code cryptosystems with high information rates. Lecture Notes in Computer Science, 434: 657–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340657.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340657.
                                        pdf.

Quisquater:1990:ZKP

[64] Jean-Jacques Quisquater and André Bouckaert. Zero-knowledge procedures for confidential access to medical records (extended summary). Lecture Notes in Computer Science, 434: 662–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340662.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340662.
                                        pdf.

DomingoiFerrer:1990:FSK

[65] Josep Domingo i Ferrer and Llorenç Huguet i Rotger. Full secure key exchange and authentication with no previously shared secrets. Lecture Notes in Computer Science, 434:665– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340665.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340665.
pdf.

Roggeman:1990:VFS

[66] Yves Roggeman. Varying feedback shift registers. Lecture Notes in Computer Science, 434:670–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340670.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340670.
pdf.

Gollmann:1990:CC

[67] Dieter Gollmann and William G. Chambers. A cryptanalysis of Step_{k,m}-cascades. Lecture Notes in Computer Science, 434:680–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340680.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340680.
pdf.

Schnorr:1990:EISa

[68] Claus P. Schnorr. Efficient identification and signatures for smart cards (abstract). Lecture Notes in Computer Science, 434:688–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/0434/04340688.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340688.
                                        pdf.

Waidner:1990:DCD

[69] Michael Waidner and Birgit Pfitzmann. The dining cryptographers in the disco: Unconditional sender and recipient untraceability with computationally secure serviceability (abstract). Lecture Notes in Computer Science, 434:690–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340690.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340690.
                                        pdf.

Carter:1990:SCL

[70] Glyn Carter. Some conditions on the linear complexity profiles of certain binary sequences. Lecture Notes in Computer Science, 434:691–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0434/04340691.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0434/04340691.
                                        pdf.

Brown:1990:DPT

[71] Lawrence Brown and Jennifer Seberry. On the design of permutation P in DES type cryptosystems. Lecture Notes in Computer Science, 434:696– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/0434/04340696.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340696.
pdf.

Agnew:1990:FEC

[72] Gordon B. Agnew, R. C. Mullin, and Scott A. Vanstone. A fast elliptic curve cryptosystem. Lecture Notes in Computer Science, 434:706– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0434/04340706.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0434/04340706.
pdf.

Anonymous:1990:AIa

[73] Anonymous. Author index. Lecture Notes in Computer Science, 434:709–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.com/
link/service/series/0558/papers/
0434/0434ind.pdf.

Kahn:1990:KGN

[74] David Kahn. Keying the German Navy’s Enigma (invited). Lecture Notes in Computer Science, 435:2– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350002.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350002.
pdf.

Desmedt:1990:MCS

[75] Yvo Desmedt. Making conditionally secure cryptosystems unconditionally abuse-free in a general context (extended abstract). Lecture Notes in Computer Science, 435:6– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350006.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350006.
pdf.

Damgaard:1990:EBC

[76] Ivan Bjerre Damgård. On the existence of bit commitment schemes and zero-knowledge proofs. Lecture Notes in Computer Science, 435:17– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350017.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350017.
pdf.

Brand:1990:PNU

[77] Russell L. Brand. Problems with the normal use of cryptography for providing security on unclassified networks (invited). Lecture Notes in Computer Science, 435:30– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350030.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350030.
pdf.

Kohl:1990:UEK

[78] John T. Kohl. The use of encryption in Kerberos for network authentication (invited). Lecture Notes in Computer Science, 435:35– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350035.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350035.
pdf.

Feldmeier:1990:UPS

[79] David C. Feldmeier and Philip R. Karn. UNIX password security — ten years later (invited). Lecture Notes in Computer Science, 435:44– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350044.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350044.
pdf.

Smith:1990:PPC

[80] Jonathan M. Smith. Practical problems with a cryptographic protection scheme (invited). Lecture Notes in Computer Science, 435:64– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350064.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350064.
pdf.

Barrett:1990:SDU

[81] Paul Barrett and Raymund Eisele. The smart diskette — A universal user token and personal crypto-engine (invited). Lecture Notes in Computer Science, 435:74–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350074.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350074.
                                        pdf.

Chan:1990:QSP

[82] Agnes Hui Chan and Richard A. Games. On the quadratic spans of periodic sequences. Lecture Notes in Computer Science, 435:82– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350082.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350082.
                                        pdf.

Jansen:1990:SFS

[83] Cees J. A. Jansen and Dick E. Boekee. The shortest feedback shift register that can generate a given sequence. Lecture Notes in Computer Science, 435:90–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350090.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350090.
                                        pdf.

Maurer:1990:PLR

[84] Ueli M. Maurer and James L. Massey. Perfect local randomness in pseudorandom sequences. Lecture Notes in Computer Science, 435:100–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350100.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350100.
pdf.

Goldreich:1990:SPD

[85] Oded Goldreich and Hugo Krawczyk. Sparse pseudorandom distributions (extended abstract). Lecture Notes in Computer Science, 435:113–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350113.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350113.
pdf.

Naor:1990:BCU

[86] Moni Naor. Bit commitment using pseudo-randomness (extended abstract). Lecture Notes in Computer Science, 435:128–??, 1990. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/0435/04350128.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0435/04350128.pdf.

Krawczyk:1990:HPC

[87] Hugo Krawczyk. How to predict congruential generators. Lecture Notes in Computer Science, 435:138–153, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350138.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350138.
pdf.

Preneel:1990:CTA

[88] Bart Preneel, Antoon Bosselaers, René Govaerts, and Joos Vandewalle. A chosen text attack on the modified cryptographic checksum algorithm of Cohen and Huang. Lecture Notes in Computer Science, 435: 154–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350154.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350154.
pdf.

Zeng:1990:LCT

[89] Kencheng Zeng, C. H. Yang, and T. R. N. Rao. On the linear consistency test (LCT) in cryptanalysis with applications. Lecture Notes in Computer Science, 435:164–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350164.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350164.
pdf.

Fiat:1990:BR

[90] Amos Fiat. Batch RSA. Lecture Notes in Computer Science, 435:175– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350175.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350175.
pdf.

Bender:1990:IEC

[91] Andreas Bender and Guy Castagnoli. On the implementation of elliptic curve cryptosystems. Lecture Notes in Computer Science, 435:186– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350186.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350186.
pdf.

Bellare:1990:NPD

[92] Mihir Bellare and Shafi Goldwasser. New paradigms for digital signatures and message authentication based on non-interactive zero knowledge proofs. Lecture Notes in Computer Science, 435:194–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350194.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350194.
pdf.

Chaum:1990:US

[93] David Chaum and Hans van Antwerpen. Undeniable signatures. Lecture Notes in Computer Science, 435: 212–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350212.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350212.
                                        pdf.

Merkle:1990:CDS

[94] Ralph C. Merkle. A certified digital signature (subtitle: That antique paper from 1979). Lecture Notes in Computer Science, 435:218– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350218.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350218.
                                        pdf.

Schnorr:1990:EISb

[95] Claus P. Schnorr. Efficient identification and signatures for smart cards. Lecture Notes in Computer Science, 435:239–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350239.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350239.
                                        pdf.

Soete:1990:SSV

[96] Marijke De Soete, Jean-Jacques Quisquater, and Klaus Vedder. A signature with shared verification scheme. Lecture Notes in Computer Science, 435:253– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350253.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350253.
pdf.

Even:1990:LLD

[97] Shimon Even, Oded Goldreich, and Silvio Micali. On-line/off-line digital signatures. Lecture Notes in Computer Science, 435:263–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/0435/04350263.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0435/04350263.pdf.

Brickell:1990:CIS

[98] Ernest F. Brickell and Daniel M. Davenport. On the classification of ideal secret sharing schemes (extended abstract). Lecture Notes in Computer Science, 435:278–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350278.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350278.
pdf.

Laih:1990:DTS

[99] Chi-Sung Laih, Lein Harn, Jau-Yien Lee, and Tzonelih Hwang. Dynamic threshold scheme based on the definition of cross-product in an Ndimensional linear space. Lecture Notes in Computer Science, 435:286– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350286.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350286.
                                        pdf.

Chor:1990:SSI

[100] Benny Chor and Eyal Kushilevitz. Secret sharing over infinite domains (extended abstract). Lecture Notes in Computer Science, 435:299–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350299.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350299.
                                        pdf.

Desmedt:1990:TC

[101] Yvo Desmedt and Yair Frankel. Threshold cryptosystems. Lecture Notes in Computer Science, 435:307–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350307.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350307.
                                        pdf.

Chick:1990:FAC

[102] Gerald C. Chick and Stafford E. Tavares. Flexible access control with master keys. Lecture Notes in Computer Science, 435:316–??, 1990. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350316.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350316.
pdf.

Tatebayashi:1990:KDP

[103] Makoto Tatebayashi, Natsume Matsuzaki, and David B. Newman, Jr. Key distribution protocol for digital mobile communication systems. Lecture Notes in Computer Science, 435: 324–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350324.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350324.
pdf.

Buchmann:1990:KES

[104] Johannes A. Buchmann and Hugh C.

Williams. A key exchange system based on real quadratic fields (extended abstract). Lecture Notes in Computer Science, 435:335–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350335.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350335.
pdf.

Yacobi:1990:KDS

[105] Yacov Yacobi and Zahava Shmuely.

On key distribution systems. Lecture Notes in Computer Science, 435: 344–??, 1990. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350344.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350344.
                                        pdf.

Nelson:1990:SAE

[106] Ruth Nelson and John Heimann. SDNS architecture and end-to-end encryption. Lecture Notes in Computer Science, 435:356–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350356.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350356.
                                        pdf.

Brickell:1990:SHI

[107] Ernest F. Brickell. A survey of hardware implementations of RSA (invited), (abstract). Lecture Notes in Computer Science, 435:368–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350368.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350368.
                                        pdf.

Findlay:1990:MEU

[108] Paul A. Findlay and Brian A. Johnson. Modular exponentiation using recursive sums of residues. Lecture Notes in Computer Science, 435:371– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/0435/04350371.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350371.
pdf.

Morita:1990:FMM

[109] Hikaru Morita. A fast modularmultiplication algorithm based on a higher radix. Lecture Notes in Computer Science, 435:387–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350387.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350387.
pdf.

Bos:1990:ACH

[110] Jurjen N. E. Bos and Matthijs J.

Coster. Addition chain heuristics. Lecture Notes in Computer Science, 435:400–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350400.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350400.
pdf.

Quisquater:1990:HECb

[111] Jean-Jacques Quisquater and Jean-Paul Delescaille. How easy is collision search. new results and applications to DES (abstract and results). Lecture Notes in Computer Science, 435:408–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/

                                        bibs/0435/04350408.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350408.
                                        pdf.

Damgaard:1990:DPH

[112] Ivan Bjerre Damgård. A design principle for hash functions. Lecture Notes in Computer Science, 435:416– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350416.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350416.
                                        pdf.

Merkle:1990:OWH

[113] Ralph C. Merkle. One way hash functions and DES. Lecture Notes in Computer Science, 435:428–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350428.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350428.
                                        pdf.

Magliveras:1990:PCP

[114] Spyros S. Magliveras and Nasir D. Memon. Properties of cryptosystem PGM. Lecture Notes in Computer Science, 435:447–??, 1990. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://link.springer-ny.com/link/
                                        service/series/0558/bibs/0435/04350447.
                                        htm;   http://link.springer-ny.
                                        com/link/service/series/0558/papers/
                                        0435/04350447.pdf.

Zheng:1990:CBC

[115] Yuliang Zheng, Tsutomu Matsumoto, and Hideki Imai. On the construction of block ciphers provably secure and not relying on any unproved hypotheses (extended abstract). Lecture Notes in Computer Science, 435: 461–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350461.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350461.
pdf.

Okamoto:1990:DZKb

[116] Tatsuaki Okamoto and Kazuo Ohta.

Disposable zero-knowledge authentications and their applications to untraceable electronic cash. Lecture Notes in Computer Science, 435:481– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350481.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350481.
pdf.

Ben-Or:1990:EIS

[117] Michael Ben-Or, Shafi Goldwasser, Joe Kilian, and Avi Wigderson. Efficient identification schemes using two prover interactive proofs. Lecture Notes in Computer Science, 435:498– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350498.htm;   http://
link.springer-ny.com/link/service/

                                        series/0558/papers/0435/04350498.
                                        pdf.

Boyar:1990:CCZ

[118] Joan Boyar and René Peralta. On the concrete complexity of zeroknowledge proofs. Lecture Notes in Computer Science, 435:507–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350507.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350507.
                                        pdf.

Feige:1990:ZKP

[119] Uriel Feige and Adi Shamir. Zero knowledge proofs of knowledge in two rounds. Lecture Notes in Computer Science, 435:526–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350526.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350526.
                                        pdf.

Kilian:1990:MRZ

[120] Joe Kilian, Silvio Micali, and Rafail Ostrovsky. Minimum resource zeroknowledge proofs (extended abstract). Lecture Notes in Computer Science, 435:545–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350545.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350545.
                                        pdf.

Bellare:1990:NIO

[121] Mihir Bellare and Silvio Micali. Noninteractive oblivious transfer and applications. Lecture Notes in Computer Science, 435:547–??, 1990. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://link.springer-ny.com/link/
service/series/0558/bibs/0435/04350547.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
0435/04350547.pdf.

Beaver:1990:MPT

[122] Donald Beaver. Multiparty protocols tolerating half faulty processors. Lecture Notes in Computer Science, 435:560–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350560.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350560.
pdf.

Cleve:1990:CGD

[123] Richard Cleve. Controlled gradual disclosure schemes for random bits and their applications. Lecture Notes in Computer Science, 435:573– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350573.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350573.
pdf.

Beaver:1990:MCF

[124] Donald Beaver and Shafi Goldwasser. Multiparty computation with faulty majority. Lecture Notes in Computer Science, 435:589–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350589.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350589.
pdf.

Chaum:1990:SDA

[125] David Chaum. The spymasters double-agent problem: Multiparty computations secure unconditionally from minorities and cryptographically from majorities. Lecture Notes in Computer Science, 435:591–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350591.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350591.
pdf.

Bellare:1990:SSK

[126] Mihir Bellare, Lenore Cowen, and Shafi Goldwasser. On the structure of secret key exchange protocols. Lecture Notes in Computer Science, 435: 604–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350604.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350604.
pdf.

Shamir:1990:EIS

[127] Adi Shamir. An efficient identification scheme based on permuted kernels (extended abstract). Lecture Notes in Computer Science, 435:606– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350606.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350606.
pdf.

Ostrovsky:1990:ESP

[128] Rafail Ostrovsky. An efficient software protection scheme (abstract). Lecture Notes in Computer Science, 435: 610–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350610.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350610.
pdf.

Adams:1990:GBE

[129] Carlisle M. Adams and Stafford E.

Tavares. Good S-boxes are easy to find. Lecture Notes in Computer Science, 435:612–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/0435/04350612.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350612.
pdf.

White:1990:CDP

[130] Steve R. White. Convert distributed processing with computer viruses. Lecture Notes in Computer Science, 435: 616–??, 1990. CODEN LNCSD9.

ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350616.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350616.
                                        pdf.

Price:1990:PDS

[131] Wyn L. Price. Progress in data security standardisation. Lecture Notes in Computer Science, 435:620– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350620.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350620.
                                        pdf.

Miyaguchi:1990:FCC

[132] Shoji Miyaguchi. The FEAL- 8 cryptosystem and a call for attack. Lecture Notes in Computer Science, 435: 624–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/0435/04350624.htm;   http://
                                        link.springer-ny.com/link/service/
                                        series/0558/papers/0435/04350624.
                                        pdf.

Quisquater:1990:HEZ

[133] Jean-Jacques Quisquater, Louis C. Guillou, and Thomas A. Berson. How to explain zero-knowledge protocols to your children. Lecture Notes in Computer Science, 435:628– ??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-

ny.com/link/service/series/0558/
bibs/0435/04350628.htm;   http://
link.springer-ny.com/link/service/
series/0558/papers/0435/04350628.
pdf.

Anonymous:1990:AIb

[134] Anonymous. Author index. Lecture Notes in Computer Science, 435:633–??, 1990. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://link.springer-ny.com/
link/service/series/0558/papers/
0435/0435ind.pdf.